Inferential.

A company offering online speed reading courses claims that students who take their courses show a 5 times (500%) increase in the number of words they can read in a minute without losing comprehension. A random sample of 100 students yielded an average increase of 415% with a standard deviation of 220%.

Calculate a 95% confidence interval for the average increase in number of words students can read in a minute without losing comprehension. Choose the closest answer.

A. (371.88, 458.12)
B.(378.7, 451.3)
C.(411.37, 418.63)
D.(412.09, 417.91)
1
Expert's answer
2020-06-08T19:52:21-0400

The critical value for "\alpha=0.05" is "z_c=z_{1-\alpha/2}=1.96".

The corresponding confidence interval is computed as shown below:


"CI=(\bar{X}-z_c\times{\sigma\over \sqrt{n}},\bar{X}+z_c\times{\sigma\over \sqrt{n}})="


"=(415-1.96\times{220\over \sqrt{100}}, 415+1.96\times{220\over \sqrt{100}})="

"=(371.88, 458.12)"

A. (371.88, 458.12)